# Software for the SpaceDREAM Robotic Arm

### v1

@@ Abstract
The software stack is based on a real time Linux as operating system and the middleware “links and nodes” providing topic communication and service calls as well as process manager functionality. The actual mission software is partitioned into low-level real time control software consisting of a hardware abstraction layer providing telemetry from the robot’s joints as well as validating and sending commands from a controller implemented in Simulink. Non-real time modules are responsible for parameterizing the controller and reading camera and other sensor data. A mission script orchestrates the individual tasks by commanding all software components. This mission script is also responsible for tracking subsystem health and potentially disabling erroneous subsystems and is itself monitored by the OBC’s hardware watchdog. Finally, a data transmission program ensures prioritized telemetry data transfer to Earth via the spacecraft. Core components of this architecture have already been tested during the 42nd DLR parabolic flight campaign under micro gravity conditions proving their effectiveness and reliability. The detailed software architecture, design choices and software tests will be described in the remainder of the paper.
@@ 1 Introduction
The SpaceDREAM project builds upon the TINA hardware [13] with the goal of a low-cost evaluation of the robotic arm and its controllers. Specifically, operations in outer space with vacuum, extreme temperature ranges as well as radiation should be tested. This paper details the software running on the On-Board Computer (OBC) which was developed for the mission under unique time and cost constraints. Section 2 details the outline of the mission to be performed ranging from simple joint-level control over compound motions with different controllers up to the integration of adaptive Virtual Fixtures (VFs) [15] which will in the future aid with teleoperated control of the robot. After analyzing the requirements in Section 3, we describe the software components that could be reused (Section 4) as well as the newly developed software (Section 5). In Section 6, we outline the integration and testing of individual components as well as of the whole system performed as of now. Section 7 concludes with an outlook to remaining steps before a mission start.
@@ 2 The SpaceDREAM Mission
Procurement of space hardware has long lead times. Hence, the principle criteria for the OBC was the availability within the project timeframe. We therefore chose the Xiphos Q8s [17] as OBC as it could be delivered within a few weeks timeframe and it combines all required features on space-grade hardware. It features a quad core ARM Cortex A53 processor for the software components and a Xilinx Zynq FPGA which allows to interface the robot’s SpaceWire communication link using an IP core as well as GPIO ports for controlling the logic and motor power of the robot as well as the cameras and LED light in the end effector. 2x of eMMC memory allow to buffer telemetry data before it is being transmitted to Earth. 4 partitions of each on two flash memories contain the operating system based on Yocto Linux which is extended by Xiphos drivers for the hardware interfaces on the FPGA and the mission software. Another important peripheral is the watchdog which is used to restart the operating system in case of non-recoverable software malfunction. Fig. 2 shows the peripherals and interfaces of the OBC connecting it to the joints of the robot, the spacecraft and other peripherals. The combination of this available OBC hardware with the mission requirements determines the software requirements which will be detailed in the next section.
@@ 3.1 General Requirements
Most critical requirement for the software development was the timeline and available resources. Initially, only 6 months were allocated for the development of a fully functional software stack followed by 3 months of testing with 3-4 full time software developers. This required a very focused development, only programming novel components where required and reusing as much software as possible. Using software with which the developers were already familiar was a key requirement speeding up programming.
@@ 3.2 Mission Execution Requirements
Most of those requirements, especially about splitting software in individual components, the communication between parts as well as the process dependency management are equivalently found on conventional robotic systems. We therefore try to match as many existing components as possible to meet these requirements while avoiding to develop new software.
@@ 4.1 Process Management
For managing processes, the links and nodes222https://gitlab.com/links_and_nodes/links_and_nodes process manager is being used. This software manages process environments by e.g. setting individual path and other environment variables. Via a configuration file, dependencies between processes can be modeled to ensure that processes start in the required order. By parsing the standard output of a process against regular expression, it can be determined whether a process is fully ready or has error output. Fig. 3 shows the gui of this process manager that can be used to inspect process execution. The same gui can also connect to the headless instance running on the OBC. In order to be able to use this process manager, small code changes were required to ensure full independency from gui libraries. As the process manager is written in Python, additionally, the Ruff linter333https://docs.astral.sh/ruff/ was used to fix potential coding issues.
@@ 4.3 Build Infrastructure
To be able to build software for robot systems with heterogeneous operating systems, the conan444https://docs.conan.io/2/ package manager is used. Recipes written in Python specify dependencies, how the software package is built from sources, packaged and they also allow to set runtime environments. Those recipes can also be used to build software for the OBC- in fact, by ensuring that the Yocto base installation contains all required system dependencies, the same dependency tree can be used. This allows to reuse well-tested dependencies of reused software components, therefore decreasing development effort and increasing the ratio of already tested software. A further element is the continuous integration infrastructure which allows to build packages on a build server for every commit and software collections at specified timepoints, which allows to always have prebuilt binaries available.
@@ 4.5 Hardware Abstraction Layer (HAL)
The SpaceWire bus used to communicate with the robot’s joints requires configuration commands to be sent at the beginning, followed by a cyclic communication. As the protocol is very similar to DLR’s David robot, the HAL developed for this robot [19] can be reused with minimal changes.
@@ 5.3 Camera Interface
In this work, we used two GoPro Hero 10 black555https://gopro.com/en/us/shop/cameras/hero10-black/CHDHX-101-master.html cameras, which are controlled via a REST API. One camera is mounted at the end effector while the other one is placed at the base of the robot (Fig. 1). As the OBC only offers one USB connection, the cameras are connected to a USB switch which needs to be switched to the camera that should be used. Once the action to take an image is called, the camera server sends a request to the camera controller, which in turn sends a request to the camera. After the shutter is triggered, the camera encodes the image and stores it on the camera’s memory card. The GoPro Hero 10 black can capture images with different fields of view but only in 4k resolution in the RGB color space with 8-bit encoding in the .jpg format. The camera controller waits for the image to be stored and then downloads the image to the OBC. At last, a post-processing step is taken to reduce the resolution and to convert the color space to the desired format.
@@ 5.4 Startup and Mission Sequence
Fig. 9 shows the startup sequence which is executed as a systemd service after booting. By default, the system should only boot once and stay active for the whole mission. In case of errors, reboots might occur which should also be handled by the software without malfunctioning of the demo or data loss. It first enables both the Ethernet as well as the SpaceWire networks and then activates the hardware Watchdog which reboots the system in case any of the following steps gets stuck. Next, one eMMC card is mounted for storing the demonstration data to be transmitted. As documented by Xiphos, the eMMC controller might be impaired by radiation effects which might render it unresponsive. It is however unlikely that the controllers for both eMMC cards fail at the same time. The startup sequence therefore selects one eMMC card at random; when the further startup fails because of a blocking controller, the Watchdog will automatically reboot the system. This random selection should then at some point select the other, working controller. In case mounting of the first selected eMMC fails, the second one is selected. If this one also cannot be mounted, a reformatting of both cards is attempted, as a last resort, a reboot is initiated to perform a reset.
@@ 5.5 Data Synchronization
As mentioned in the requirements, we assume to only have a downlink but no response channel through which acknowledgments of transmission can be sent. The data synchronization software therefore needs to work without acknowledgement, it therefore resends data depending on the importance to compensate for potential packet losses and transmission errors. The most important data is sent first to ensure it is being transferred before a potential deorbit of the spacecraft. The protocol is designed to transmit data using UDP and a low transfer rate of around , however, using TCP, the spacecraft could apply backpressure to communicate the available bandwidth. Checksums then allow to verify the integrity of the transferred data.
For transferring data, files are split in fragments containing a checksum. The packet size to be transferred depends on the underlying network’s capabilities, each packet consists of one or multiple fragments. Depending on the connection quality, a higher or lower fragment size can be chosen - smaller fragments are beneficial with frequent transmission errors while bigger fragment sizes reduce the overhead coming from the transmission of checksums. Three parameters allow to configure the transmission of fragments coming from a file: a priority value, a resend number and a minimum waiting time between retransmission. Based on those parameters, fragments are ordered in a priority queue for sending. On the receiver side, those file fragments are merged again. Missing data leads to holes, i.e. zero-filled fragments in the data. This necessites the usage of robust file formats - e.g. using the jpeg format for images is beneficial as missing fragments do not corrupt the whole image. As the image header does not have this robustness property, it is duplicated on file level when recording images to have a higher chance of retrieving the image metadata.
@@ 6 Integration and Testing
With all reused and new software components ready, the overall mission software can be put together. Fig. 11 shows an overview of all those elements. Except for the data synchronization process, all software components are started from the links and nodes manager. The first process to be started is the digital joint power which enables power for the digital electronics of the robotic arm. This is a prerequesite for SpaceWire communication and must therefore be started before starting the HAL. Before enabling power, the process first disables power to achieve a reset of the robot’s joints - it also does not communicate with the other processes as the GPIO needs to be active during the whole mission. The other processes communicate using links and nodes topics, services, and parameters. The lnrecorder process records all communication at a specified rate; this data is then sent to Earth via the data synchronization process for later analyis.
@@ 6.1 Code Quality
Even before testing the code, a set of linters was applied to the whole newly generated code base to ensure a good quality even though it was not possible to write unit tests for every function. To this end, cppcheck was applied to all C++ code to find potential issues already during development. For Python code, a combination of the Ruff linter and mypy for type checking was applied - all new Python code was fully typed to avoid issues from type errors. The black formatter and isort were furthermore used to avoid whitespace code changes when different developers were working on the same code.
@@ 6.2 Simulated Development
As explained in Section 3, no hardware was available at the beginning of the software implementation, also, new implementations should be tested in simulation first in order not to damage the EM. To therefore enable a simulation environment for development, a small dynamic simulation of the robot was set up using Simulink. Combined with a visualization and camera simulation in RViz (Fig. 12), the whole software stack as implemented on the OBC can be run on a desktop computer as well. This simulation environment is set up in a custom configuration for the links and nodes manager which also contains the data synchronization process to also be able to test the data transfer with the actual amount of data written.
@@ 6.5 Pending Tests
While the HAL and controller model could be tested extensively during the parabolic flight, tests for other central parts were not yet conducted. In fact, during development, an unforseeable delay of the flight became apparent which halted programming of specialized components. Therefore, the full implementation of the mission sequence is yet to be completed from the experience during the parabolic flights. This implementation also depends on the exact available workspace and acceleration limits on the spacecraft which have not been communicated to date. Once this information is available, a test of the full mission should be conducted to ensure that it meets workspace and power requirements and that all data can be transferred successfully from the OBC to a simulated ground station with injected defects.

### v2

@@ Abstract
The software stack is based on a real time Linux as operating system and the middleware ”links and nodes” providing topic communication and service calls as well as process manager functionality. The actual mission software is partitioned into low-level real time control software consisting of a hardware abstraction layer providing telemetry from the robot’s joints as well as validating and sending commands from a controller implemented in Simulink. Non-real time modules are responsible for parameterizing the controller and reading camera and other sensor data. A mission script orchestrates the individual tasks by commanding all software components. This mission script is also responsible for tracking subsystem health and potentially disabling erroneous subsystems and is itself monitored by the OBC’s hardware watchdog. Finally, a data transmission program ensures prioritized telemetry data transfer to Earth via the spacecraft. Core components of this architecture have already been tested during the 42nd DLR parabolic flight campaign under micro gravity conditions proving their effectiveness and reliability. The detailed software architecture, design choices and software tests will be described in the remainder of the paper.
@@ 1 Introduction
The SpaceDREAM project builds upon the TINA hardware [13] with the goal of a low-cost evaluation of the robotic arm and its controllers. Specifically, operations in outer space with vacuum, extreme temperature ranges as well as radiation should be tested. This paper details the software running on the On-Board Computer (OBC) which was developed for the mission under unique time and cost constraints. While existing frameworks already tested in space such as DLR Outpost [15], NASA’s F Prime [16, 17] or core Flight System [18] could have been used to develop the flight software, we have instead opted to reuse as much software from DLR’s ground robots as possible. Main reason for this decision was the familiarity of the developers with existing software and the potential to reuse many existing software components that have already proven their reliability through years of operation on different robotic systems. A re-implementation of e.g. already available control software for other frameworks in the required quality level would have been impossible within the project’s timeframe.
The outline of the mission to be performed is detailed in Section 2, ranging from simple joint-level control over compound motions with different controllers up to the integration of adaptive Virtual Fixtures (VFs) [19] which will in the future aid with teleoperated control of the robot. After analyzing the requirements in Section 3, we describe the software components that could be reused (Section 4) as well as the newly developed software (Section 5). In Section 6, we outline the integration and testing of individual components as well as of the whole system performed as of now. Section 7 concludes with an outlook to remaining steps before a mission start.
@@ 2 The SpaceDREAM Mission
Procurement of space hardware has long lead times. Hence, the principle criteria for the OBC selection was its availability within the project timeframe. We therefore chose the Xiphos Q8s [21] as OBC as it could be delivered timely and it combines all required features on space-grade hardware. It features a quad core ARM Cortex A53 processor for the software components and a Xilinx Zynq FPGA which allows to interface the robot’s SpaceWire communication link using an IP core as well as GPIO ports for controlling the logic and motor power of the robot as well as the cameras and LED light in the end effector. 2x of eMMC memory allow to buffer telemetry data before it is being transmitted to Earth. 4 partitions of each on two flash memories contain the operating system based on Yocto Linux which is extended by Xiphos drivers for the hardware interfaces on the FPGA and the mission software. Another important peripheral is the watchdog which is used to restart the operating system in case of non-recoverable software malfunction. Fig. 2 shows the peripherals and interfaces of the OBC connecting it to the joints of the robot, the spacecraft and other peripherals. The combination of this available OBC hardware with the mission requirements determines the software requirements which will be detailed in the next section.
@@ 3.1 General Requirements
Most critical requirement for the software development was the timeline and available resources. Initially, only 6 months were allocated for the development of a fully functional software stack followed by 3 months of testing with 3-4 full time software developers. This required a very focused development, only programming novel components where required and reusing as much software as possible. Using software with which the developers were already familiar was a key requirement to speed up programming.
@@ 3.2 Mission Execution Requirements
Most of those requirements, especially about splitting software in individual components, the communication between those parts as well as the process dependency management are equivalently found on conventional robotic systems. We therefore try to match as many existing components as possible to meet these requirements while avoiding to develop new software.
@@ 4.1 Process Management
For managing processes, the links and nodes222https://gitlab.com/links_and_nodes/links_and_nodes process manager is being used. This software manages process environments by e.g. setting individual path and other environment variables. Via a configuration file, dependencies between processes can be modeled to ensure that processes start in the required order. By parsing the standard output of a process against regular expressions, it can be determined whether a process is fully ready or has error output. Fig. 3 shows the gui of this process manager that can be used to inspect process execution. The same gui can also connect to the headless instance running on the OBC. In order to be able to use this process manager, small code changes were required to ensure full independency from gui libraries. As the process manager is written in Python, additionally, the Ruff linter333https://docs.astral.sh/ruff/ was used to fix potential coding issues.
@@ 4.3 Build Infrastructure
To be able to build software for robot systems with heterogeneous operating systems, the conan444https://docs.conan.io/2/ package manager is used. Recipes written in Python specify dependencies, how the software package is built from sources, packaged and they also allow to set runtime environment variables. Those recipes can also be used to build software for the OBC- in fact, by ensuring that the Yocto base installation contains all required system dependencies, the same dependency tree can be used. This allows to reuse well-tested dependencies of reused software components, therefore decreasing development effort and increasing the ratio of already tested software. A further element is the continuous integration infrastructure which allows to build packages on a build server for every commit and software collections at specified timepoints, which allows to always have prebuilt binaries available.
@@ 4.5 Hardware Abstraction Layer (HAL)
The SpaceWire protocol used to communicate with the joints of the robot requires configuration commands to be sent at the beginning, followed by a cyclic communication. As the protocol is very similar to DLR’s David robot, its HAL [23] can be reused with minimal changes.
@@ 5.3 Camera Interface
In this work, we used two GoPro Hero 10 black555https://gopro.com/en/us/shop/cameras/hero10-black/CHDHX-101-master.html cameras, which are controlled via a REST API. One camera is mounted at the end effector while the other one is placed at the base of the robot (Fig. 1). As the OBC only offers one USB connection, the cameras are connected to a USB switch which needs to be switched to the camera that should be used. Once the action to take an image is called, the camera server sends a request to the camera controller, which in turn sends a request to the camera. After the shutter is triggered, the camera encodes the image and stores it on the camera’s memory card. The GoPro Hero 10 black can capture images with different fields of view but only in 4k resolution in the RGB color space with 8-bit encoding in the .jpg format. The camera controller waits for the image to be stored and then downloads the image to the OBC. Finally, a post-processing step reduces the resolution and converts the color space to the desired format.
@@ 5.4 Startup and Mission Sequence
Fig. 9 shows the startup sequence which is executed as a systemd service after booting. By default, the system should only boot once and stay active for the whole mission. In case of errors, reboots might occur which should also be handled by the software without malfunctioning of the demo or data loss. It first enables both the Ethernet as well as the SpaceWire networks and then activates the hardware Watchdog which reboots the system in case any of the following steps gets stuck. Next, one eMMC card is mounted for storing the demonstration data to be transmitted. As documented by Xiphos, the eMMC controller might be impaired by radiation effects which might render it unresponsive. It is however unlikely that the controllers for both eMMC cards fail at the same time. The startup sequence therefore selects one eMMC card at random; when the further startup fails because of a blocking controller, the Watchdog will automatically reboot the system. This random selection then at some point selects the other, working controller. In case mounting of the first selected eMMC fails, the second one is selected. If this one also cannot be mounted, a reformatting of both cards is attempted, as a last resort, a reboot is initiated.
@@ 5.5 Data Synchronization
As mentioned in the requirements, we assume to only have a downlink but no response channel through which acknowledgments of transmission can be sent. The data synchronization software therefore needs to work without acknowledgement, it thus resends data depending on the importance to compensate for potential packet losses and transmission errors. The most important data is sent first to ensure it is being transferred before a potential deorbit of the spacecraft. The protocol is designed to transmit data using UDP and a low transfer rate of around , however, using TCP, the spacecraft could apply backpressure to communicate the available bandwidth. Checksums then allow to verify the integrity of the transferred data.
For transferring data, files are split in fragments containing a checksum. The packet size to be transferred depends on the underlying network’s capabilities, each packet consists of one or multiple fragments. Depending on the connection quality, a higher or lower fragment size can be chosen - smaller fragments are beneficial with frequent transmission errors while bigger fragment sizes reduce the overhead coming from the transmission of checksums. Three parameters allow to configure the transmission of fragments coming from a file: a priority value, a resend number and a minimum waiting time between retransmission. Based on those parameters, fragments are ordered in a priority queue for sending. On the receiver side, those file fragments are merged again. Missing data leads to holes, i.e. zero-filled fragments in the data. This necessites the usage of robust file formats - e.g. using the jpeg format for images is beneficial as missing fragments do not corrupt the whole image. As the image header does not have this robustness property, it is duplicated on file level to have a higher chance of successful metadata transmission.
@@ 6 Integration and Testing
With all reused and new software components ready, the overall mission software can be put together. Fig. 11 shows an overview of all those elements. Except for the data synchronization process, all software components are started from the links and nodes manager. The first process to be started is the digital joint power which enables power for the digital electronics of the robotic arm. This is a prerequesite for SpaceWire communication and must therefore be started before starting the HAL. Before enabling power, the process first disables power to achieve a reset of the robot’s joints - it also does not communicate with the other processes as the power needs to be active during the whole mission. The other processes communicate using links and nodes topics, services, and parameters. The lnrecorder process records all communication at a specified rate; this data is then sent to Earth via the data synchronization process for later analyis.
@@ 6.1 Code Quality
Even before testing the code, a set of linters was applied to all new code base to ensure a good quality even though it was not possible to write unit tests for every function. To this end, cppcheck was applied to all C++ code to find potential issues already during development. For Python code, a combination of the Ruff linter and mypy for type checking was applied - all new Python code was fully typed to avoid issues from type errors. The black formatter and isort were furthermore used to avoid whitespace code changes when different developers were working on the same code.
@@ 6.2 Simulated Development
As explained in Section 3, no hardware was available at the beginning of the software implementation, also, new implementations should be tested in simulation first in order not to damage the EM. To therefore enable a simulation environment for development, a small dynamic simulation of the robot was set up using Simulink. Combined with a visualization and camera simulation in RViz (Fig. 12), the whole software stack as implemented on the OBC can be run on a desktop computer as well. This simulation environment is set up in a custom configuration for the links and nodes manager which also contains the data synchronization process to also be able to test the data transfer with the actual amount of data.
@@ 6.4 Parabolic Flight Testing
Both the controller as well as the HAL process are part of the real time control loop of the robot. While no specific load measurements on the OBC were performed, real time execution at a rate of could be ensured during the whole flight with an average jitter of in the controller. As the data was stored locally, all messages and service calls were recorded leading to a data rate of . Assuming an available downlink of , this data rate has to be reduced for a space flight, which is easily possible by e.g. reducing the recording rate of most topics.
@@ 6.5 Pending Tests
While the HAL and controller model could be tested extensively during the parabolic flight, tests for other central parts were not yet conducted. In fact, during development, an unforseeable delay of the flight became apparent which halted programming of specialized components. Therefore, a full implementation of the mission sequence with the exact robotic movements is yet to be completed from the experience during the parabolic flights. This implementation also depends on the available workspace for the robot as well as acceleration limits on the spacecraft which have not been communicated to date.
Once this information is available, a finalized implementation of the mission sequence is possible. This full mission should be tested to ensure that it meets physical workspace, OBC load and power requirements and that all data can be transferred successfully from the OBC to a simulated ground station with injected defects using the expected data rate. Key components to be tested are furthermore the startup sequence (Fig. 9) with injected eMMC faults as well as the automatic mission execution (Fig. 10) which have not yet been tested thoroughly as they were not required for the parabolic flight. The same applies for the camera and its software.
Regular testing will also allow to calibrate nominal sensor measurement values which can be used during the robot checkout performed at the start of the mission sequence (Section 2). Finally, tests of the only once executed HDRM release sequence as well as restarting the robot from arbitrary configurations after an OBC reboot have to be tested. With testing those individual components as well as the full mission, the whole system will be ready for a space flight.
@@ Acknowledgements.
- [15] DLR-RY, “Dlr outpost.” [Online]. Available: https://github.com/DLR-RY/outpost-core
- [16] R. Bocchino, T. Canham, G. Watney, L. Reder, and J. Levison, “F prime: an open-source framework for small-scale flight software systems,” 2018.
- [17] T. Canham, “The mars ingenuity helicopter - a victory for open-source software,” in 2022 IEEE Aerospace Conference (AERO). IEEE, pp. 01–11.
- [18] D. McComas, J. Wilmot, and A. Cudmore, “The core flight system (cfs) community: Providing low cost solutions for small spacecraft,” in Annual AIAA/USU conference on small satellites, no. GSFC-E-DAA-TN33786, 2016.